Two capacitors are connected in series c1=1f , v1=10v and c2=3f , v2= 20f find total voltage of the system 30 is not the answer
1
Expert's answer
2020-02-11T13:44:38-0500

Two capacitors are connected in series:


"q_1=C_1V_1=(1)(10)=10\ C""q_2=C_2V_2=(3)(20)=60\ C"

Total charge:


"q=q_1+q_2=10+60=70\ C"

Capacitance:


"C=\frac{C_1C_2}{C_1+C_2}=\frac{(1)(3)}{1+3}=0.75\ F"

Total voltage of the system:


"U=\frac{q}{C}=\frac{70}{0.75}=93.
3\ V"